For the data 1, 2, 2, 3, 3, 4, 4, 4, 4, . . . . . 9, 9, the product of mean and mode equals please advice method to solve this.

Grade:10

1 Answers

Harsh Patodia IIT Roorkee
8 years ago
As you can see any number is occuring the number of time its value is.

So 9 is clearly the mode in given numbers
For mean you can add like ( 1x1 + 2x2 + 3x3 + . . . . 9x9) = \sum_{n=1}^{9} n2
= n(n+1)(2n+1)/6
=285

Total no of observations is ( 1+2 + 3. . . 9) = \sum_{n=1}^{9} n = n(n+1)/2 = 45
Mean = 285/45 = 19/3

Product= 19*9/3 = 57